Random sample of size 4 are drawn from the finite population which consists of the 

numbers 2,3,7,8,and 10.



1
Expert's answer
2021-03-08T13:25:17-0500

a. the mean of the population


"\mu=\dfrac{2+3+7+8+10}{5}=6"

the standard deviation of the population


"\sigma^2=\dfrac{(2-6)^2+(3-6)^2+(7-6)^2+(8-6)^2+(10-6)^2}{5}"

"=9.2"

"\sigma=\sqrt{9.2}=3.033"


b. There are "\dbinom{5}{4}" samples of size two which can be drawn without replacement: 


"\dbinom{5}{4}=\dfrac{5!}{4!(5-4)!}=5"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample\ No. & Sample\ value & Sample\ mean\ \bar{X} \\ \hline
 1 & (2,3,7,8) & 5 \\
 \hdashline
 2 & (2,3,7,10) & 5.5 \\
 \hdashline
 3 & (2,3,8,10) & 5.75 \\
 \hdashline
 4 & (2,7,8,10) & 6.75 \\
 \hdashline
 5 & (3,7,8,10) & 7 
\end{array}"

The sampling distribution of the sample mean "\bar{X}" and its mean and standard deviation are:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 5 & 1 & 0.2 & 1 & 5 \\
 \hdashline
 5.5 & 1 & 0.2 & 1.1 & 6.05 \\
 \hdashline
 5.75 & 1 & 0.2 & 1.15 & 6.6125 \\
 \hdashline
 6.75 & 1 & 0.2 & 1.35 & 9.1125 \\
 \hdashline
 7 & 1 & 0.2 & 1.4 & 9.8 \\
 \hdashline
 Total & 5 & 1 & 6 & 36.575
\end{array}"

"E(\bar{X})=\sum\bar{X}f(\bar{X})=6"


"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(E(\bar{X}))^2=36.575-6^2"

"=0.575"

Verification:


"\mu=6=E(\bar{X})"

"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})="

"=\dfrac{9.2}{4}(\dfrac{5-4}{5-1})=0.575"